OpenStudy (anonymous):

Which pair of angles has congruent values for the sin x° and the cos y°? (5 points) Select one: a. 18°; 18° b. 18°; 162 c. 18°; 72° d. 18°; 108°

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

is it c?

OpenStudy (science0229):

I got c, too

OpenStudy (anonymous):

sin 18 = cos 72, right?

OpenStudy (anonymous):

so what is it

OpenStudy (science0229):

You can see this easily if you draw a right triangle.

OpenStudy (anonymous):

18 and 72 are congruent so is 20 and 70 30 and 60 and so on... can u see the pattern ? 90 - a = b , a and b are congruent

OpenStudy (anonymous):

So is the Answer C

OpenStudy (science0229):



OpenStudy (science0229):

Compare sin18 and cos 72 to discover that they're the same.

OpenStudy (science0229):

And as @kx2bay said, \[\sin(x)=\cos(90-x)\]or\[\cos(x)=\sin(90-x)\] in degrees

OpenStudy (anonymous):

man i got a warning from the moderators for giving out straight answers and got told to lead u to it insetad, here me and science are trying our best to stick to code of conduct and help u ;)

OpenStudy (anonymous):

are we still getting C for the answer

OpenStudy (anonymous):

if u got C then u r right :)

OpenStudy (anonymous):

Thnx guys

OpenStudy (anonymous):

cheers :)